Hi really need your help on this question.
Question: Within a rectangular garden 10m wide and 20 long, we wish to have a walk around the borders of uniform width so as to leave an area of 96m^2 for flowers. How wide should be?

Answer by Boreal(15235)

the uniform area around has x m width.
The length of the flower bed will be 20-2x
The width is 10-2x
This equals 96 m^2
200-60x+4x^2=96
rewrite as
4x^2-60x+104=0
factor out a 4
4(x^2-15x+26)=0
4(x-13)(x-2)=0
x=13,2. The first can't work.
It is a 2 meter border.
the length for flowers is 16 m (20-2(2))
the width for flowers is 6 m
96m^2